# Towards Goal-oriented Large Language Model Prompting: A Survey

### v1

@@ 2.5 Selecting Valuable Sub-goals for Effective Goal Achievement
In the first stage where we decompose a goal into a sub-goal sequence, most existing works rely on LLM’s knowledge to finish the process. Due to the hallucination problem, it is difficult to guarantee that all of the decomposed sub-goals are correct and relevant. One of the solutions, described in Section 2.1, is to reduce sub-goal space before asking LLMs to decide sub-goals. We denote such methods as sub-goal preprocess. Alternatively, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions. Then, we only select those valuable sub-goals from candidates to pursue a global optimal goal achievement. We denote this selection process as sub-goal postprocess. In this section, we introduce methods for sub-goal postprocess, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al. (2023) and its variants, and finally Graph of thoughts (GoT) Besta et al. (2023). An illustration of these approaches is shown in Fig. 2.
@@
As a naive way of prompting LLMs, CoT asks LLMs to decompose the goal by themselves. The sub-goal selection is implicitly incorporated into the autoregressive decoding process where tokens with higher probabilities are more likely to be generated. However, the selected sub-goals may be irrelevant or even wrong due to the hallucination problem, which leads to the failure of goal achievement.
Four works are proposed to improve the robustness of CoT through sub-goal postprocess. Self-consistency Wang et al. (2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of the final output to decide valuable sub-goal sequences. The result from the chosen sequence is considered the final result. MCR Yoran et al. (2023) improves Self-consistency by considering all of the sub-goals in various sequences instead of the sub-goal sequence result. After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and take the result from the meta path. Recmind Wang et al. (2023b) and Selection-Inference Creswell et al. (2022) expand the sub-goal to multiple candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
@@
ToT Yao et al. (2023) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure. To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search or depth-first search. For the comparison of sub-goals, ToT solely relies on the self-evaluation of LLMs. CoT and its variants can be considered special cases of ToT.
RAP Hao et al. (2023b) and GDP-Zero Yu et al. (2023) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS). Compared with ToT, MCTS has an additional step for backpropagation, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence. This operation strikes a proper balance between exploration and exploitation to find valuable sub-goals efficiently, which makes MCTS outperform heuristic-based search algorithms on complex or less structured tasks.
@@
GoT Besta et al. (2023) advances ToT by modeling all of the LLM-generated sub-goals as a graph. While in ToT, only sub-goals in a sequence or at the same step can interact with each other through comparison, lookahead, or backtracking, GoT allows the synergy of arbitrary sub-goals. Specifically, GoT supports sub-goal aggregation to generate a new sub-goal and sub-goal refinement based on other dependent sub-goals. Finally, valuable sub-goals are selected via LLM self-evaluation on the constructed graph.
3 Application
Application Sub-category Stage
Method
Reasoning Arithmetic Reasoning ①②③④⑤
CoT Wei et al. (2022), Self-consistency Wang et al. (2022), Least-to-most Zhou et al. (2022), PoT Chen et al. (2022), Self-refine Madaan et al. (2023), Plan-and-solve Wang et al. (2023a), MWP Zhang et al. (2023), RAP Hao et al. (2023b), SelfCheck Miao et al. (2023), Toolformer Schick et al. (2023), Successive Prompt Dua et al. (2022)
Commonsense Reasoning
CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023a), Toolformer Schick et al. (2023)
Symbolic Reasoning
Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023a)
Logical Reasoning
Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023b)
Planning Virtual Environment ①②④
Zero-shot Planner Huang et al. (2022a), Re-prompting Raman et al. (2022), SALP Gramopadhye and Szafir (2023), RAP Hao et al. (2023b), Reflexion Shinn et al. (2023), LLM+P Liu et al. (2023a), SayPlan Rana et al. (2023), DEPS Wang et al. (2023c), GITM Zhu et al. (2023), Inner Monologue Huang et al. (2022b)
Real Environment
Inner Monologue Huang et al. (2022b)
Question Answering Multihop QA ①③④⑤
DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
Open domain QA
Toolformer Schick et al. (2023)
Code Generation
Reflexion Shinn et al. (2023), Self-debug Chen et al. (2023), Self-refine Madaan et al. (2023)
Dialogue ①④⑤
ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
Recommendation
Recmind Wang et al. (2023b)
Table 2: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1. We maintain a list of related datasets and miscellaneous tasks at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering.
In this section, we provide an overview of the application of goal-oriented prompt engineering on LLMs. Specifically, we first introduce the definition of the task and then describe how current methods, following into different stages within our framework, leverage goal-oriented prompt engineering on LLMs to achieve task objectives. A summary of the applicable tasks, stages involved, and respective research works are shown in Table 2. Note that here we only introduce representative tasks. For miscellaneous tasks and related datasets, we maintain a list at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering to give more comprehensive information. We can see that our framework can be applied to a wide range of tasks.
Reasoning.
There are specifically four types of reasoning tasks: arithmetic reasoning, commonsense reasoning, logical reasoning, and symbolic reasoning. These reasoning tasks focus on math word problems, world knowledge, logical principles and rules, and manipulation of symbols such as letters, respectively. Reasoning involves multiple intermediate steps to obtain a final answer. Additionally, to obtain the correct answer the LLM must interpret the question correctly, extract necessary information from the given context, and maintain consistency within its intermediate steps.
The goal of these tasks can be considered as producing an answer to a given question. The intermediate steps can be viewed as sub-goals that are sequentially connected to reach the final goal Wei et al. (2022); Wang et al. (2023a); Khot et al. (2022). The sub-goal results from action implementation, which in this case are intermediate conclusions, are often evaluated and revised to avoid error accumulation which may negatively influence the following derivations Madaan et al. (2023); Miao et al. (2023). Alternatively, for certain actions like calculations, we could leverage an external calculator to ensure a precise sub-goal result Schick et al. (2023) and employ sketches to help equation formulation (action selection) Zhang et al. (2023). We can also ask LLMs to explore parallel possible intermediate conclusions to form an overall better solution.
Planning.
Given an objective like “Fetch a bottle from the kitchen”, planning involves developing a series of sub-objectives (e.g. “Go to the kitchen”, “Find the cupboard”) Liu et al. (2023a); Song et al. (2023) and reaching sub-objectives by selecting appropriate actions (e.g. “Move”, “Search”) Hao et al. (2023a); Raman et al. (2022) to achieve the overall goal in a virtual or real environment.
We can see that works focusing on this task do not propose methods for action implementation and valuable sub-goal selection. This is because the evaluated environments are relatively straightforward and existing works mainly focus on the accuracy of planning without considering efficiency. LLMs often interact with virtual environments through APIs Wang et al. (2023c); Zhu et al. (2023) or real environments through mechanical devices Huang et al. (2022b) to ensure the actions are implemented precisely.
Question Answering.
Open domain question answering is a task aimed at answering factual questions without any explicit evidence. Multihop QA is an advanced task over open-domain QA. The former relies more on world knowledge while the latter requires multi-step reasoning based on world knowledge to answer more complex questions.
Considering answering complex questions as the goal, LLMs are asked to first decompose them into basic sub-questions Khot et al. (2022); Sun et al. (2023), gain knowledge for sub-questions Schick et al. (2023), and reflect on the correctness of the knowledge Press et al. (2022); Shinn et al. (2023). Sometimes LLMs may ask irrelevant sub-questions or give wrong answers to those sub-questions. To avoid their negative influence, there are methods proposed to lead LLMs to select valuable sub-questions from all candidates Yoran et al. (2023).
Code Generation.
Code generation aims to write a code snippet according to the given description. In this task, our goal is to generate code that is both functionally correct and aligns with the description. All of the works applied to this task fall into the evaluation of results. They leverage the error messages from unit tests Shinn et al. (2023); Chen et al. (2023) or rely on LLMs themselves Chen et al. (2023); Madaan et al. (2023) to reflect the validity of generated code snippets and revise them if needed.
Dialogue.
Dialogue systems are designed to provide functional service or social support to users via interactions. LLMs often play the role of a conversational agent to achieve certain goals (e.g. query clarification, offer emotional support). To give users positive interaction, LLMs have to correctly understand the context and generate clear and brief responses. Thus, some sub-goals are pre-defined according to the goal to ensure effective decomposition Deng et al. (2023). And LLMs are asked to reflect on their response to generate better content Madaan et al. (2023).
Recommendation.
The recommender system, which targets providing personalized recommendations to users based on their preferences, is crucial in many application scenarios including social media, e-commerce websites, etc. Considering giving ratings for items as the final goal, collecting rating information from various sources are then sub-goals. To rely on more valuable sources, LLMs can be asked to compare their credibility and authenticity, and finally take the most valuable source as the overall ratings Wang et al. (2023b).
@@ 4 Challenges & Opportunities
In this section, we discuss the challenges and opportunities of goal-oriented prompt engineering. Even if there have been numerous works proposed to facilitate processing at each stage, there remain some potential directions.
Synergy of stages.
The synergy of each stage in our framework will further improve the performance of goal-oriented prompt engineering. We find that though numerous works are proposed to address different stages in the framework, no one synergizes them all.
For reasoning tasks, each stage has specially proposed methods, and putting them together is straightforward. However, for some tasks like code generation, the vacancy of works for certain stages hinders the integration. According to Table 2, there are only works focusing on sub-goal evaluation in code generation. To bridge the gap, we have to understand how the code generation process can fit into our framework. In fact, programming is a process of breaking down a problem into small, manageable parts and solving them in a logical sequence. For example, suppose we are tasked with calculating the average of a set of numbers. We can break this problem down into the following steps: (1) Input a set of numbers. (2) Calculate the sum of these numbers. (3) Determine the count of numbers. (4) Use the sum and count to calculate the average. Considering these intermediate steps as sub-goals, we decompose the goal of finishing a programming problem into a sequence of sub-goals. For other stages like action selection and sub-goal evaluation, we could apply methods that help LLMs efficiently select actions to calculate the summation or reflect whether their intermediate answer is correct.
Application to other tasks.
Though existing works have covered up to 11 tasks, goal-oriented prompt engineering has the potential for broader applications. For example, apart from implementing functions, code review also plays a crucial role in software development. Various review criteria can be taken as sub-goals to ensure the quality of the code, which is our overall goal. We may prompt LLMs to design approaches to evaluate different criteria (e.g. text-based similarity measurement for coding style evaluation, unit tests generation for functionality verification), which corresponds to the action selection to achieve sub-goals. Similar to Program-of-thoughts Chen et al. (2022), we can use external interpreters to test the code. For some criteria without precise success or failure signals, we can apply sub-goal self-evaluation methods to facilitate better feedback. This fitting-in is also suitable for requirement analysis where we also have to think through several perspectives. Taking one step further, we think this framework even has the potential to be applied to the whole life cycle of software development.
@@
The works we review in this paper mainly address the accuracy of goal-oriented prompt engineering. In other words, they care whether the goal is achieved finally. However, efficiency also plays a crucial role in application. As we are optimistic about the accuracy when employing methods in each stage of our framework to prompt design, we think efficiency is the next key factor that affects performance. For example, in planning tasks, LLMs are expected to form a plan that costs the least energy of human beings or robots. And in dialogue systems, users prefer to achieve their goals with as few interaction rounds as possible. To evaluate efficiency, we need to either design heuristic rules for automatic measurement or annotate the resource-consuming information by human beings.
Hierarchical decomposition.
Existing works decompose the goal into a sub-goal sequence, which can be limiting. The method might fall short in adequately addressing the intricacies of each sub-goal, especially in situations where each sub-goal itself is complex and multifaceted. Thus, LLMs may not achieve certain sub-goals perfectly, then the error will be accumulated, leading to the failure of overall goal achievement. To make sure each sub-goal is simple enough to be handled by LLMs, we introduce the idea of hierarchical decomposition where sub-goals can be decomposed into even more simple sub-goals. GITM Zhu et al. (2023) could be viewed as employing the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones. Though a significant success rate increase by employing hierarchical decomposition is reported, relying on solely a tree may not be able to handle more complex problems, which leaves room for further improvement.

### v2

@@ 2.5 Further Selection of Valuable Sub-goals
In the first stage where we decompose a goal into a sub-goal sequence, most existing works rely on LLM’s knowledge to finish the process. Due to the hallucination problem, it is difficult to guarantee that all of the decomposed sub-goals are correct and relevant. To address this problem, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions. Then, we only select those valuable sub-goals from candidates to pursue a global optimal goal achievement. In this section, we introduce methods for such sub-goal postprocessing, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al. (2023) and its variants, and finally Graph of thoughts (GoT) Besta et al. (2023). An illustration of these approaches is shown in Fig. 2 to better distinguish these methods.
@@
We consider CoT as simply selecting all sub-goals since LLMs have access to all of the sub-goals during the generation process, which makes it prone to irrelevant sub-goals or wrong sub-goal results. Four works are proposed to improve the robustness of CoT through sub-goal postprocess. Self-consistency Wang et al. (2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of different outputs from each COT sequence to decide valuable sub-goal sequences. The result from the chosen sequence is considered the final result. MCR Yoran et al. (2023) improves Self-consistency by selecting and combining sub-goals from different sequences. After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and predict the result based on the meta path. Recmind Wang et al. (2023c) and Selection-Inference Creswell et al. (2022) allow multiple sub-goal candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
@@
ToT Yao et al. (2023) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure. To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search. For the comparison of sub-goals, ToT solely relies on the self-evaluation of LLMs. CoT and its variants can be considered special cases of ToT, where each node only has one child node.
RAP Hao et al. (2023) and GDP-Zero Yu et al. (2023) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS). Compared with ToT, MCTS has an additional back-propagation step, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence. This operation strikes a proper balance between exploration and exploitation to find valuable sub-goals efficiently, which makes MCTS outperform heuristic-based search algorithms on complex or less structured tasks.
@@
GoT Besta et al. (2023) advances ToT by modeling all of the LLM-generated sub-goals as a graph. While in ToT, only sub-goals in a sequence or at the same step can interact with each other through comparison, lookahead, or backtracking, GoT allows the synergy of arbitrary sub-goals. What’s more, GoT supports sub-goal aggregation to generate a new sub-goal and sub-goal refinement based on other dependent sub-goals. Finally, valuable sub-goals are selected via LLM self-evaluation on the constructed graph.
3 Discussion
Application.
After understanding how each method leverages goal-oriented prompt engineering to achieve goals with LLMs, we find it insightful to provide a bird-eye view of the surveyed paper. In Table 9, we classify these papers by their targeted tasks and the stages involved in our goal-oriented framework, to help readers better understand the current progress [^1] Note that here we only introduce representative tasks. For more comprehensive information, we maintain a list at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering..
The 36 representative works in our survey cover 11 tasks, including arithmetic reasoning (11 works), commonsense reasoning (5 works), symbolic reasoning (2 works), logical reasoning (3 works), planning in virtual environment (10 works), planning in real environment (1 work), multihop question answering (5 works), open-domain question answering (1 work), code generation (4 works), dialogue systems (3 works), and recommender systems (1 work). We could see that 60% of the works focus on reasoning, 31% of the works are for planning and the remaining 9% of the works are for other tasks.
For reasoning tasks, all the surveyed works involve only a single stage, with goal decomposition, action execution, sub-goal result evaluation, and valuable sub-goal selection holding 23% equally. There is potential for exploring the synergy of stages, which we will discuss in Section 4. For planning tasks, goal decomposition shares 31%, action selection shares 23%, and sub-goal result evaluation shares the rest 46%. No surveyed work for planning tasks involves the action execution stage and valuable sub-goal selection stage. This is because the currently available evaluation benchmarks are based on either simple real environments or carefully designed virtual environments. Actions are either easy or can be assured to be perfectly executed. Nonetheless, as the field develops, we can expect more complicated environments when actions cost differently or cannot be perfectly executed. Under such scenarios, there should be attention paid to action execution and valuable sub-goal selection stage.
For those minority tasks like code generation and recommendation, most of the stages remain unexplored. With the goal-oriented framework proven effective in reasoning and planning tasks, it should be valuable to explore the potential of the framework on other tasks and we will discuss it in Sec 4.
Performance.
Furthermore, We compare the performance of the surveyed methods in Appendix A (Table 2, 3, 4, 5, 6, 7, 8). Compared with vanilla CoT, on average goal decomposition (Plan-and-solve) improves by 6.1%, action selection (PoT) improves by 2.9%, and valuable sub-goal selection (Self-consistency) improves by 22.7% in arithmetic reasoning, as shown in Table 6. Besides, we observe a gradual increase in performance when integrating additional stages into the method. When combined with Self-consistency, Plan-and-solve can be improved by 11.3% and PoT can be improved by 5.2% on average across several arithmetic reasoning datasets, which indicates that stage synergy may bring further improvement.
During data collection, we find it hard to compare between methods due to the variety of LLMs used, the number of in-context examples included, and the evaluation metric. With our survey, we hope to give researchers a better estimate of the current progress of goal-oriented prompt engineering in different tasks and also promote a unified evaluation procedure for each task for fair comparison of future methods.
@@ 4 Challenges & Opportunities
In this section, we discuss the challenges and opportunities of goal-oriented prompt engineering based on our findings.
Hierarchical Goal Decomposition.
Existing works only decompose the goal into one layer of sub-goals. However, the sub-goal itself can still be complex and multifaceted. When a certain sub-goal is complex, LLMs may not achieve it perfectly, and the error will propagated, leading to the failure of the final goal. To make sure each sub-goal is simple enough to be handled by LLMs, researchers are now exploring hierarchical decomposition, which further decomposes sub-goals into simpler "sub-sub-goals". Li et al. (2023) is a pioneering work that constructs a dataset for the script generation task focusing on hierarchical decomposition. GITM Zhu et al. (2023) employs the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones. Though a significant success rate increase by employing hierarchical decomposition is reported, relying solely on one tree may not be able to handle more complex problems, which leaves room for the introduction of graphs for further improvement.
@@
The current design of goal-oriented prompt engineering focuses on improving the accuracy of models. Nevertheless, efficiency also plays a crucial role in the real-world application of the methods. Researchers should consider the trade-off between accuracy and efficiency when necessary.
We propose to consider two perspectives of efficiency measurement, task-agnostic efficiency and task-specific efficiency. Task-agnostic efficiency aims at measuring the efficiency of model targets for any task, for instance, the total number of prompting times before reaching the sub-goal or final goal. On the other hand, task-specific efficiency is designed to measure efficiency that appears only in certain tasks.
As an example, the number of iterations that a certain method prompts LLMs would affect the speed of plan generation in planning tasks. Such task-agnostic efficiency is important in scenarios like autonomous driving path planning when we need a quick response for better user experiences. On the other hand, the output plan needs to fulfill other criteria, such as minimizing traveling time or distance. Here, the time and distance are task-specific efficiency.
Stage Synergy.
Overall there is a lack of systematic synergy of stages in any task. We found two works that combine their methods, on one stage, with a method, on another stage, in their ablation studies and show performance improvement Chen et al. (2022); Wang et al. (2023b). This suggests the power of synergizing the stages in the goal-oriented framework and careful integration of stages would have the potential for further improvement.
We take the arithmetic reasoning task as an example. To solve a math problem, a naive solution would be to follow Plan-and-solve Wang et al. (2023b) to decompose the math problem into a few logical steps. Then, following MWP Zhang et al. (2023), an operation prediction module can be applied to determine the correct mathematical operations. LLMs can select the operands for performing the operation. Then, LLMs could call an external calculator to calculate the equation to ensure an accurate result. With the result, we could follow Reflexion Shinn et al. (2023) to ask LLMs to double-check the intermediate steps again. We could further repeat the described process to get multiple sequences of sub-goals and follow MCR Yoran et al. (2023) to select valuable sub-goals for more accurate answers. Of course, a naive combination of methods from all stages is heavy, requiring additional attention paid to efficiency.
Additionally, it is also important to consider the between-stages interaction. For the same math problem, if we apply a calculator for calculation (action execution) to guarantee a correct answer, we should be careful when designing the result evaluation stage. If we use LLMs for result evaluation, the correct result from the calculator may inversely be considered as wrong due to the hallucination problem.
Broader Applications.
From our summarization, we observe an absence of study in certain stages for some tasks, such as code generation and recommender system. There is potential to improve current methods by studying the unexplored stages.
In code generation, all of the surveyed methods focus on leveraging sub-goal evaluation to refine the code generation result. Additional refinement may be done by incorporating other stages. Given a description of the desired code snippets, we could first decompose it into several required sub-functions, and then select appropriate algorithms to realize the functionality of the sub-function. As for the implementation stage of these sub-functions, we could search relevant documents of potential well-written function APIs for reference instead of generating the code fully based on LLMs’ knowledge. APIs in external libraries are regularly updated which can perform better than the static knowledge stored in LLMs.
In recommender systems, works have been proposed for enhancing action execution and sub-goal result evaluation. There is an absence of work on action selection. Take keeping user retention as the overall goal, one sub-goal might be to make sure that users are recommended with their preferred items. There are various actions we could take to predict the preference of a user for a certain item, such as inferring based on the click rate of similar items and inferring based on the watching time of relevant advertisements. Careful design on dynamically selecting between available actions based on the characteristics of the user may enable more accurate prediction.
Moreover, goal-oriented prompt engineering has the potential for broader applications, in for example software development and medical diagnosis, beyond the 11 tasks summarized in this paper.
For instance, code review is a crucial but cumbersome task in software development. It ensures the quality and maintainability of code but requires a reviewer to spend a lot of time doing the systematic examination. Automatic code review can help save such significant effort. Due to the coding ability shown by recent LLMs, there is potential to surpass existing static analysis tools in automatic code review if equipped with our goal-oriented framework. We found one primary study Tufano et al. (2024), which focuses on action selection, already shows that reducing the action space by identifying the type of request changes can significantly improve the success rate of ChatGPT. Here we give another possibility which covers more stages of our framework. With reviewing code as the goal, sub-goals can be documentation quality, function reusability, code style consistency, variable naming readability, and so on. Taking variable naming readability as the sub-goal, we can have LLMs select action between checking the self-consistency of the variable names or comparing those names with the pre-defined naming convention (the action selection stage). Subsequently, LLMs can execute the action by either calling an external tool for similarity measurement or following given heuristic rules (the action execution stage). After going through all the sub-goals (i.e. reviewing criteria), LLMs can revise the given code based on reviews from each perspective.
5 Conclusion
In this paper, we provide a review of existing research on LLMs prompting from a goal-oriented perspective. We explicitly summarize the task-agnostic internal steps involved in designing and refining prompting procedures into five stages, classify the surveyed methods by their targeted tasks, and provide a comprehensive analysis on task coverage and performance. Though significant progress has been made in this field, further research is needed to fully unleash the potential by studying hierarchical goal decomposition, the synergy of stages, application to other tasks, and considering efficiency problems. We hope this paper can serve as an overview of the current state of goal-oriented prompt engineering and stimulate further research on this important topic.
Appendix A Peroformance Comparison
In this section, we compare the performance of surveyed methods under several tasks. Due to the different settings in papers, we cannot directly take the reported metric score. To address this problem, we first collect scores from those published papers. Note that we not only take the score of the proposed methods in those papers but also that of baselines if they appear as one of our surveyed works. Once the settings of the same method in different papers have no overlap with each other, we take the score under the setting that has the most similar performance. For preprint papers, we take their result once their compared baselines are already recorded and the discrepancy is no more than 3%. As a result, the reported performance in our survey is not strictly a fair comparison, and we don’t make conclusions stating which method works best in certain contexts. We mainly aim to show that the introduction of goal-oriented prompt design can consistently bring improvement. The results are shown in Table , 2, 3, 4, 5, 6, 7, 8.
Methods VirtualHome
Zero-shot Planner 46.1
Re-prompting 41.8
SALP 49.0
Table 2: Performance comparison on planning in virtual environment. The evaluation metric is solve rate (%).
Metric Methods CSQA StrategyQA
Solve rate CoT 79.0 75.3
Self-consistency 81.5 81.6
Plan-and-solve 71.9 65.4
F1 MCR - 73.6
Self-consistency - 72.2
Self-ask - 69.3
Table 3: Performance comparison on commonsense reasoning under solve rate (%) and F1 (%).
Methods HumanEval MBPP
CoT 66.5 48.8
Self-refine 65.2 48.8
Reflexion 91.0 77.1
Self-debug - 75.6
INTERVENOR 75.6 69.8
Table 4: Performance comparison on code generation. The evaluation metric is Pass@1 (%).
Methods Amazon Review Yelp
CoT 0.7167 0.9794
ToT 0.7059 0.9766
Selection-Inference 0.6892 0.9698
Table 5: Performance comparison on rating prediction in recommender system. The evaluation metric is Mean Absolute Error.
Methods GSM8K SVAMP ASDiv AQuA AddSub MultiArith
CoT 56.4 69.9 74.0 38.9 85.3 83.8
Self-consistency 74.4 86.6 81.9 48.3 93.7 99.3
PoT 57.0 70.8 - 43.9 85.1 92.2
Plan-and-solve 59.3 75.7 - 46.0 92.2 91.8
Table 6: Performance comparison on arithmetic reasoning. The evaluation metric is solve rate (%).
Methods Symbolic Reasoning Logical Reasoning
last letter concatenation coinflip PrOntoQA
CoT 70.4 99.0 87.8
Self-consistency 73.4 99.5 89.8
Plan-and-solve 75.2 99.6 -
RAP - - 94.2
Table 7: Performance comparison on symbolic reasoning and logical reasoning. The evaluation metric is solve rate (%).
Metric Methods 2WikiMultihopQA MuSiQue HotpotQA
Accuracy CoT 29.8 12.6 -
Least-to-most 29.0 16.8 -
Self-ask 40.1 16.3 -
F1 MCR 67.9 - 57.0
Self-consistency 65.4 - 56.4
Self-ask 63.8 - 50.2
DecomP 64.1 29.3 56.4
Table 8: Performance comparison on multihop question answering under solve rate (%) and F1 (%).
Application Sub-category Stage Method
Reasoning Arithmetic Reasoning ①②③④⑤ CoT Wei et al. (2022), Self-consistency Wang et al. (2022), Least-to-most Zhou et al. (2022), PoT Chen et al. (2022), Self-refine Madaan et al. (2023), Plan-and-solve Wang et al. (2023b), MWP Zhang et al. (2023), RAP Hao et al. (2023), SelfCheck Miao et al. (2023), Toolformer Schick et al. (2023), Successive Prompt Dua et al. (2022)
Commonsense Reasoning CoT Wei et al. (2022), Self-consistency Wang et al. (2022), MCR Yoran et al. (2023), Plan-and-solve Wang et al. (2023b), Toolformer Schick et al. (2023)
Symbolic Reasoning Self-consistency Wang et al. (2022), Plan-and-solve Wang et al. (2023b)
Logical Reasoning Selection-Inference Creswell et al. (2022), DecomP Khot et al. (2022), RAP Hao et al. (2023)
Planning Virtual Environment ①②④ Zero-shot Planner Huang et al. (2022a), Re-prompting Raman et al. (2022), SALP Gramopadhye and Szafir (2023), RAP Hao et al. (2023), Reflexion Shinn et al. (2023), LLM+P Liu et al. (2023a), SayPlan Rana et al. (2023), DEPS Wang et al. (2023d), GITM Zhu et al. (2023), Inner Monologue Huang et al. (2022b)
Real Environment Inner Monologue Huang et al. (2022b)
Question Answering Multihop QA ①③④⑤ DecomP Khot et al. (2022), Self-ask Press et al. (2022), Reflexion Shinn et al. (2023), MCR Yoran et al. (2023), PEARL Sun et al. (2023)
Open domain QA Toolformer Schick et al. (2023)
Code Generation Reflexion Shinn et al. (2023), Self-debug Chen et al. (2023), Self-refine Madaan et al. (2023), INTERVENOR Wang et al. (2023a)
Dialogue ①④⑤ ProCoT Deng et al. (2023), GDP-Zero Yu et al. (2023), Self-refine Madaan et al. (2023)
Recommendation ③⑤ Recmind Wang et al. (2023c)
Table 9: Applicable tasks from existing works. The stage column indicates the stages involved in the corresponding methods with regard to the goal-oriented framework in Fig. 1.